OpenStudy (maheshmeghwal9):

If \[z, \omega \ne 0.\] & \[|z|= |\omega|.\] and \[\arg.(z)+\arg.(\omega)=\pi \space \space then \space \space z=...?\]

OpenStudy (maheshmeghwal9):

Answer is \[- \bar \omega\]

OpenStudy (experimentx):

seems that they are conjugate of each other. what is 'w' BTW??

OpenStudy (maheshmeghwal9):

omega

OpenStudy (experimentx):

i know ... i mean what is it's value?? like is it root of unity or something??

OpenStudy (experimentx):

or does it have some specific value??

OpenStudy (maheshmeghwal9):

cube root of unity

OpenStudy (maheshmeghwal9):

if not specified we take cube root of unity.

OpenStudy (experimentx):

\[ \omega^3 = 1 = e^{2 \pi} \implies \omega = \cos \left ( 2 \pi \over 3 \right ) - \sin \left ( 2 \pi \over 3 \right )\]

OpenStudy (experimentx):

z is it's conjugate.

OpenStudy (maheshmeghwal9):

how we come to know that "z is it's conjugate." ?

OpenStudy (maheshmeghwal9):

i think there should be \[ \omega = \cos \left ( 2 \pi \over 3 \right ) - i \sin \left ( 2 \pi \over 3 \right )\]

OpenStudy (maheshmeghwal9):

i think a typing mistake:)

OpenStudy (experimentx):

Nop ... + sorry i did not pay attention.

OpenStudy (maheshmeghwal9):

but np:)

OpenStudy (experimentx):

the argument of w is 2pi/3 the sum of argument is pi ... so find the argument of z

OpenStudy (maheshmeghwal9):

pi/3

OpenStudy (experimentx):

so z = cos pi/3 + i sin pi/3

OpenStudy (maheshmeghwal9):

ya i think too:)

OpenStudy (experimentx):

the modulus of z should be 1 .. so only value you can have is the above value.

OpenStudy (maheshmeghwal9):

Isn't there any logical & stepwise method?????

OpenStudy (experimentx):

i don't know.

OpenStudy (maheshmeghwal9):

np:) but thanx a lot for help:)